OpenStudy (anonymous):

What is the value of 3 to the power 2 over 3 to the power 4

OpenStudy (anonymous):

Welcome to OpenStudy. The rule for dividing powers with the same base is to SUBTRACT the exponents. For example,\[\frac{ 4^5 }{ 4^2 } = 4^{5-2} = 4^3\]Try it with your problem. What do you get for an answer?